# The Evolution of Dataset Distillation: Toward Scalable and Generalizable Solutions

### v1

@@ 1 Introduction
TABLE I: Comparison of Recent Dataset Distillation Surveys
@@ 2.1.4 Latent and Frequency Space Methods
Complementary to latent space approaches, frequency-domain methods offer another efficient avenue for dataset distillation. Shin et al. [41] introduced Frequency Domain-based Dataset Distillation (FreD), a method that operates in the frequency domain rather than the spatial domain. FreD optimizes a selective subset of important frequency dimensions based on explained variance ratios, avoiding the need to process the full dimensionality of the spatial domain. This approach drastically lowers the memory budget per synthetic instance while preserving critical information. Yang et al. [42] introduced Neural Spectral Decomposition (NSD), leveraging the low-rank structure of datasets to decompose them into spectral tensors and kernel matrices, enhancing learning efficiency and integrating seamlessly with frameworks like MTT.
@@ 2.2.1.2   Diffusion-based Methods
where $T$ denotes the total number of diffusion steps, $t$ represents the current timestep in the backward process, $\mu_{\theta}$ is the parameterized mean predictor network with learnable parameters $\theta$, and ${z}_{T}$ is the initial noisy state. The linear interpolation coefficients $(1-\frac{t}{T})$ and $\frac{t}{T}$ control the balance between the mean prediction and initial state, providing enhanced gradient flow during training without compromising the generation quality. This framework demonstrates superior performance for high-resolution tasks while integrating seamlessly with existing algorithms. Taking a different approach to latent space optimization, Qin et al. [65] proposed Distributional Dataset Distillation (D3) and applied it into the federated learning setting [66]. In their method, D3 represents classes as Gaussian distributions in latent space, optimizing the parameters to generate images using a decoder network. Collectively, these latent space methods emphasize the power of leveraging compact and expressive representations, offering scalable, efficient, and high-quality solutions for dataset distillation across diverse architectures and tasks.
@@ 2.4.1 Addressing Non-IID Challenges
Self-supervised and Transfer Learning While most dataset distillation research has focused on supervised learning, self-supervised and transfer learning [74, 113, 112] remain underexplored. Lee et al. proposed KRR-ST [119], which uses kernel ridge regression and mean squared error objectives to remove randomness from the distillation process. This approach facilitates efficient pretraining with unlabeled data, significantly reducing computational costs. Further, Joshi et al. introduced MKDT [120], a two-stage method that stabilizes self-supervised learning (SSL) trajectories by first training student models via knowledge distillation from SSL-trained teachers. In alignment with these stabilized trajectories, MKDT achieves substantial performance gains in SSL pretraining and downstream tasks.
@@ 2.5.1 Adversarial Attack
Figure 5: Overview of DD-RobustBench [140]. Image from [140].
@@ 3 Emerging Applications and Domains
3.1 Temporal Domain: Video and Audio
@@ 3.3 Medical Domain
3.4 Scientific Analysis
Dataset distillation has demonstrated its versatility in tackling scientific challenges, such as galaxy morphology analysis [165], and practical applications, like object detection [166], image super-resolution [167]. Guan et al. [165] applied dataset distillation to galaxy morphology analysis, proposing Self-Adaptive Trajectory Matching (STM) to improve upon MTT. STM uses statistical hypothesis testing to adaptively monitor validation loss, ensuring efficient stopping criteria and reduced hyperparameter tuning. Their work includes a curated high-confidence version of the Galaxy Zoo 2 dataset [168], successfully distilling morphological features and achieving superior convergence and stability. For object detection, Qi et al. [166] introduced DCOD, the first dataset condensation framework for this task, delivering competitive results on benchmarks like Pascal VOC [169] and MS COCO [170]. Dietz et al. [167] demonstrate the first comprehensive exploration of dataset distillation for image super-resolution (SR) [171], successfully reducing the original dataset size by 91.12% while maintaining comparable model performance through both pixel-space and latent-space distillation techniques, thereby offering a promising memory- and computation-efficient approach.
Dataset distillation has also shown promise in mitigating noise-related challenges. Cheng et al. [172] proposed leveraging dataset distillation as a novel denoising tool for learning with noisy labels, addressing challenges like feedback loops in traditional noise evaluation strategies. Through extensive experiments with methods like DATM, DANCE, and RCIG, they demonstrate the effectiveness of dataset distillation in removing random and natural noise, though it struggles with structured asymmetric noise. This approach also improves training efficiency and privacy by avoiding direct use of noisy data and enabling offline processing.
@@ 4.2 Performance Analysis Across Dataset Scales and Complexities
Methods Schemes Venue CIFAR-10 CIFAR-100 Tiny ImageNet ImageNet-1K
@@ 4.3 Performance and Scalability Across Techniques
Recent methods have also explored importance-based sampling and diversity-driven approaches. Notable examples include AutoPalette [86], which achieves 58.6% accuracy on CIFAR-10 at IPC=1, and EDC [76], which demonstrates strong performance across different scales, reaching 87.0% on CIFAR-10 at IPC=50. These innovations suggest that combining multiple strategies, such as diversity enhancement, importance sampling, and efficient optimization, may be key for both scalability and performance.
In summary, while traditional matching-based methods remain effective for smaller datasets, the field has evolved towards more sophisticated approaches that combine de-coupling mechanisms like SRe2L, diversity-enhancing strategies, and novel paradigms such as latent and frequency-based methods. To advance the scalability and robustness of dataset distillation, future research must continue to address the challenges of large-scale datasets, particularly focusing on methods that can maintain high performance while scaling to datasets like ImageNet-21K.
@@ 6 Conclusion
Several key trends have emerged during this period. First, the development of more sophisticated optimization strategies, such as SRe2L frameworks and diversity-driven approaches, has enabled effective distillation of large-scale datasets like ImageNet-1K and ImageNet-21K. Second, the introduction of soft labels and decoupling mechanisms has significantly improved both efficiency and performance. Third, the emergence of generative models, particularly diffusion-based approaches, has opened new possibilities for high-quality synthetic data generation. Despite these advances, important challenges remain. The scalability to ultra-large datasets, the trade-off between compression and performance, and cross-architecture generalization continue to be active areas of research. Additionally, the extension to multi-modal and temporal domains presents both opportunities and challenges that require innovative solutions.
We hope that this comprehensive survey of recent dataset distillation advances will serve as a valuable resource for researchers and practitioners, providing insights into the latest progress, existing challenges, and promising future directions in this rapidly evolving field.

### v2

@@ 1 Introduction
TABLE I: Comparison of recent dataset distillation surveys. Unlike previous surveys, which primarily focus on small datasets, low IPC settings, and supervised classification tasks, our work expands the scope to large-scale datasets, high IPC settings, generative approaches, decoupling-based solutions, and broader applications, such as self-supervised learning and federated learning.
@@ 2.1.4 Latent and Frequency Space Methods
Complementary to latent space approaches, frequency-domain methods offer another efficient avenue for dataset distillation. Shin et al. [41] introduced Frequency Domain-based Dataset Distillation (FreD), a method that operates in the frequency domain rather than the spatial domain. FreD optimizes a selective subset of important frequency dimensions based on explained variance ratios, avoiding the need to process the full dimensionality of the spatial domain. This approach drastically lowers the memory budget per synthetic instance while preserving critical information. Building upon frequency-domain techniques, Yang et al. [42] introduced Neural Spectral Decomposition (NSD), leveraging the low-rank structure of datasets to decompose them into spectral tensors and kernel matrices, enhancing learning efficiency and integrating seamlessly with frameworks like MTT. Moving beyond traditional representations, Shin et al. [43] introduces Distilling Dataset into Neural Field (DDiF), a novel parameterization framework that compresses large datasets into synthetic neural fields under limited storage budgets, offering superior performance on various datasets.
@@ 2.2.1.2   Diffusion-based Methods
where $T$ denotes the total number of diffusion steps, $t$ represents the current timestep in the backward process, $\mu_{\theta}$ is the parameterized mean predictor network with learnable parameters $\theta$, and ${z}_{T}$ is the initial noisy state. The linear interpolation coefficients $(1-\frac{t}{T})$ and $\frac{t}{T}$ control the balance between the mean prediction and initial state, providing enhanced gradient flow during training without compromising the generation quality. This framework demonstrates superior performance for high-resolution tasks while integrating seamlessly with existing algorithms. Taking a different approach to latent space optimization, Qin et al. [66] proposed Distributional Dataset Distillation (D3) and applied it into the federated learning setting [67]. In their method, D3 represents classes as Gaussian distributions in latent space, optimizing the parameters to generate images using a decoder network. To guide the distilled data generation process, in [68], Chen et al. introduced Influence-Guided Diffusion (IGD) that uses a latent diffusion model based on a Diffusion Transformer [69] to generate synthetic training datasets. They designed an influence guidance function to steer diffusions towards generating training-effective data and a deviation guidance function to enhance diversity, enabling the method to significantly improve performance across various datasets, particularly on high-resolution ImageNet-1K benchmarks.
@@ 2.4.1 Addressing Non-IID Challenges
Self-supervised and Transfer Learning While most dataset distillation research has focused on supervised learning, self-supervised and transfer learning [76, 115, 114] remain underexplored. Lee et al. [121] proposed KRR-ST, which employs kernel ridge regression and mean squared error objectives to reduce randomness in the distillation process. This method facilitates efficient pretraining with unlabeled data, significantly lowering computational costs. Expanding on this, Yu et al. [122] introduced an enhanced self-supervised DD framework that builds upon KRR-ST through three key innovations: image and representation parameterization, predefined augmentation, and approximation networks. These refinements improve cross-architecture generalization, transfer learning performance, and storage efficiency. Furhter, Joshi et al. introduced MKDT [123], a two-stage method that stabilizes SSL trajectories by first training student models via knowledge distillation from SSL-trained teachers. In alignment with these stabilized trajectories, MKDT achieves substantial performance gains in SSL pretraining and downstream tasks.
@@ 2.5.1 Adversarial Attack
Figure 5: Overview of DD-RobustBench. Image from [143].
@@ 3 Emerging Applications and Domains
3.1 Temporal Domain
@@
Sequential Data Zhang et al. [160] introduce Tucker Decomposition based Dataset Distillation (TD3), a novel Tucker Decomposition-based framework for dataset distillation in sequential recommendation, demonstrating an innovative approach to generating compact yet expressive synthetic sequence summaries.
@@ 3.3 Medical Domain
3.4 Other Applications
Scientific Discovery Dataset distillation has demonstrated its versatility in tackling scientific challenges, such as galaxy morphology analysis [169]. Guan et al. [169] applied dataset distillation to galaxy morphology analysis, proposing Self-Adaptive Trajectory Matching (STM) to improve upon MTT. STM uses statistical hypothesis testing to adaptively monitor validation loss, ensuring efficient stopping criteria and reduced hyperparameter tuning. Their work includes a curated high-confidence version of the Galaxy Zoo 2 dataset [170], successfully distilling morphological features and achieving superior performance.
Object Detection and Super-Resolution For object detection, Qi et al. [171] introduced DCOD, the first dataset condensation framework for this task, delivering competitive results on benchmarks like Pascal VOC [172] and MS COCO [173]. Dietz et al. [174] demonstrate the first comprehensive exploration of dataset distillation for image super-resolution (SR) [175], successfully reducing the original dataset size by 91.12% while maintaining comparable model performance through both pixel-space and latent-space distillation techniques, thereby offering a promising memory- and computation-efficient approach.
Enhancing Data Quality Dataset distillation has shown promise in noise-related challenges and improving dataset quality. Cheng et al. [176] proposed leveraging dataset distillation as a novel denoising tool for learning with noisy labels, addressing challenges like feedback loops in traditional noise evaluation strategies. Through extensive experiments with methods like DATM, DANCE, and RCIG, they demonstrate the effectiveness of dataset distillation in removing random and natural noise, though it struggles with structured asymmetric noise. This approach also improves training efficiency and privacy by avoiding direct use of noisy data and enabling offline processing. Building upon the insights into noisy label handling through distillation, Wu et al. [177] proposed Trust-Aware Diversion (TAD). Rather than just using distillation as a denoising tool [176], TAD introduces a dual-loop optimization framework that actively manages noisy labels during the distillation process itself. In TAD, an outer loop separates data into trusted and untrusted spaces to guide distillation, while an inner loop recalibrates untrusted samples to maximize their utility. UniDetox [178] utilized dataset distillation approach to universally mitigate toxic content in large language models by systematically refining training datasets, thereby reducing harmful biases while preserving model performance across different domains.
@@ 4.2 Performance Analysis Across Dataset Scales and Complexities
1=Methods Schemes Venue CIFAR-10 CIFAR-100 Tiny ImageNet ImageNet-1K
@@ 4.3 Performance and Scalability Across Techniques
Recent methods have also explored importance-based sampling and diversity-driven approaches. Notable examples include AutoPalette [88], which achieves 58.6% accuracy on CIFAR-10 at IPC=1, and EDC [78], which demonstrates strong performance across different scales, reaching 87.0% on CIFAR-10 at IPC=50. These innovations suggest that combining multiple strategies, such as diversity enhancement, importance sampling, and efficient optimization, may be key for both scalability and performance. In summary, while traditional matching-based methods remain effective for smaller datasets, the field has evolved towards more sophisticated approaches that combine de-coupling mechanisms like SRe2L, diversity-enhancing strategies, and novel paradigms such as latent and frequency-based methods. To advance the scalability and robustness of dataset distillation, future research must continue to address the challenges of large-scale datasets, particularly focusing on methods that can maintain high performance while scaling to datasets like ImageNet-21K.
@@ 5.1 Challenges
Fair Evaluation of Dataset Distillation Methods Recent DD methods are shifting from hard labels to soft labels to enhance model accuracy [54]. However, soft labels inherently transfer additional knowledge from a teacher model, making it difficult to isolate the true informativeness of distilled data [179, 55]. Additionally, the use of diverse data augmentation techniques during evaluation further skews results, making it unclear whether performance gains stem from the quality of the distilled data or external enhancements [179]. Furthermore, different dataset distillation methods employ varied evaluation strategies, such as distinct loss functions, leading to inconsistencies in performance comparison.
@@ 5.2 Future Directions
Enhancing Fair Evaluation and Comparisons In this survey, we present two comparative tables summarizing key aspects of various dataset distillation methods, including their use of soft labels, architectures, and other relevant factors, providing researchers with a reference to determine the most suitable approach for their needs. Additionally, Li et al. introduced DD-Ranking [179], a benchmark designed to decouple the effects of knowledge distillation and data augmentation, ensuring a fair evaluation framework to assess the true informativeness of distilled data. By combining structured comparisons with insights from existing benchmarks such as DD-Ranking [179], we aim to help researchers navigate the landscape of dataset distillation methods while promoting more standardized evaluation practices.
@@ 6 Conclusion
Several key trends have emerged during this period. First, the development of more sophisticated optimization strategies, such as SRe2L frameworks and diversity-driven approaches, has enabled effective distillation of large-scale datasets like ImageNet-1K and ImageNet-21K. Second, the introduction of soft labels and decoupling mechanisms has significantly improved both efficiency and performance. Third, the emergence of generative models, particularly diffusion-based approaches, has opened new possibilities for high-quality synthetic data generation. Despite these advances, important challenges remain. The scalability to ultra-large datasets, the trade-off between compression and performance, and cross-architecture generalization continue to be active areas of research. Additionally, the extension to multi-modal and temporal domains presents both opportunities and challenges that require innovative solutions. We hope that this comprehensive survey of recent dataset distillation advances will serve as a valuable resource for researchers and practitioners, providing insights into the latest progress, existing challenges, and promising future directions in this rapidly evolving field.
Acknowledgment
The authors would like to express their gratitude to the contributors of the Awesome-Dataset-Distillation repository [^1] https://github.com/Guang000/Awesome-Dataset-Distillation, which has served as a valuable resource in compiling recent advancements.